OpenStudy (anonymous):

PLEASE HELP ME Which value for x makes the sentence true? 2x - 3 = 3 2 4 3 0

OpenStudy (hysusonic):

plug in the numbers for x

OpenStudy (anonymous):

0

OpenStudy (texaschic101):

2x - 3 = 3 2x = 3 + 3 2x = 6 x = 6/2 x = 3 check.. 2(3) - 3 = 3 6 - 3 = 3 3 = 3 (correct) x = 3 makes it true

OpenStudy (anonymous):

last one

OpenStudy (dtan5457):

3. 2(3)=6 6-3=3.

OpenStudy (anonymous):

3

OpenStudy (hysusonic):

guys let the asker do their own work

OpenStudy (anonymous):

thanks